Solve the equation 3x^2 -6x = -3.
Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12844)

3x^2 -6x = -3
3x^2-6x+3=0
3(x^2-2x+1)=0
(x-1)^2=0
x-1=0
x=1 ans.

Answer by stanbon(75887)

Solve the equation 3x^2 -6x = -3
3x^2-6x+3 = 0
Factor:
3(x^2-2x+1) = 0
Factor:
3(x-1)^2 = 0
Solve: x-1 = 0
x = 1
==================
Cheers,
Stan H.